### v2

# Quadratic Mode Couplings in Rotating Black Holes and Their Detectability

Abstract
Quadratic quasinormal modes encode fundamental properties of black hole spacetimes. They are also one of the key ingredients of nonlinearities of General Relativity in the ringdown stage of binary black hole coalescence. In this work, we classify all possible quadratic coupling channels of quasinormal modes for a generic Kerr black hole, and use a frequency-domain pseudospectral code with hyperboloidal slicing to calculate these couplings. After accounting for all the channels in systems with reflection symmetry, our results become consistent with those extracted from numerical simulations and time-domain fits. We explore several observationally relevant quadratic modes and their detectability with third-generation ground-based detectors like Cosmic Explorer, as well as the space-borne detector LISA (Laser Interferometer Space Antenna). We find that several quadratic modes are detectable by both detectors. In addition, we make a theoretical prediction for how the signal-to-noise ratio of a quadratic quasinormal mode depends on the properties of progenitor binaries, offering insight into selecting golden events in future ringdown data analyses.
I Introduction
Black hole (BH) ringdown after binary BH coalescence encodes key information about the spacetime of the final BH [1, 2, 3], as well as the binary’s inspiral properties, such as the mass ratio, spins, etc [4, 5, 6, 7, 8, 9]. Theoretically, it is important to characterize various linear and non-linear signals in the ringdown stage and their connection to the inspiral parameters, as predicted by General Relativity (GR). In particular, understanding nonlinear signals not only enables direct tests of nonlinearities in GR, but also allows more theoretically meaningful ringdown fits at earlier times. This is particularly important for data analysis, as fitting the ringdown from earlier times generally leads to a higher signal-to-noise ratio (SNR).
Among various BH non-linearities [10, 11, 12, 13, 14, 7, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], quadratic quasinormal modes (QQNMs) stand out due to their conceptual simplicity and their potential observability using current ringdown fitting techniques [27, 28, 29, 30, 31, 32, 33]. QQNMs have been identified by fitting numerical time-domain simulation data at infinity [10, 11, 12, 13, 14, 7] and on the BH horizon [17]. However, a complete and accurate characterization of the modes relies on frequency domain methods, which is the focus of this work.
Along with an earlier work [18] (hereafter Paper I), we develop two independent frequency-domain methods to handle QQNMs for generic Kerr BHs. In Paper I, we analytically reconstructed the metric and solved the frequency-domain second-order Campanelli–Lousto-Teukolsky equation [34] along a complex contour. Alternatively, in this paper, we adopt the reduced second-order Teukolsky equation [35, 36] and use the pseudo-spectral method to numerically assemble its source term. Subsequently, we solve the equation spectrally in hyperboloidal slicing [37, 38, 39]. Both methods generate consistent results for the coupling coefficient of the prograde mode $\ell=m=2,n=0$ for BH spins $\chi\in[0,0.95]$. Furthermore, we extend our discussions in Paper I by studying more observationally relevant QQNMs, and by providing a detailed classification of all possible mode-coupling channels. After examining mode coupling through different channels and symmetry conditions in (some of the) previous numerical simulations, we find consistency with the time-domain fits for mode amplitudes.
Leveraging the recent ringdown surrogate model [^1] See also [9]. [8], we obtain theoretical predictions for how the amplitudes of quadratic modes depend on the properties of progenitor binaries. Assuming the sensitivities of Cosmic Explorer (CE) and Laser Interferometer Space Antenna (LISA) [40, 41], we carry out a parameter survey to forecast the SNRs of several leading quadratic modes in non-precessing binaries, complementing the recent study by Yi et al. [42]. Our findings indicate that several QQNMs are within the detectable range of both detectors.
Throughout this work we adopt the geometric unit by setting $G=c=1$. Complex conjugates are represented by overlines. We label a QNM with four indices $(l,m,n,p)$, where $p={\rm sgn}({\rm Re}(\omega))$. QNMs come in pairs. Their frequency values are related via
$\omega^{p}_{\ell mn}=-\bar{\omega}^{-p}_{\ell-mn},$ (1)
which indicates that the modes with $(m,p)$ and $(-m,-p)$ (and the same $l,n$) are mirrors of each other.
II QQNMs with hyperboloidal slicing
QNMs are known to diverge at spatial infinity and at the bifurcate horizon. While this divergence poses no physical issue because they do not lie in the region where the QNMs are a good approximation, it can cause numerical challenges. In Paper I, we deal with the divergence by using the complex plane approach. Here, we adopt an alternative method — hyperboloidal time slicing — which avoids the problematic regions. This enables direct numerical treatments of the reduced second-order Teukolsky equation [35, 36]
$\mathcal{T}\left[\Psi_{4}^{(2)}\right]=S_{4}^{(2)}.$ (2)
Using the hyperboloidal slicing, it has been shown that QNMs are eigenvectors of the non-normal time evolution operator [43]. The separability of the Teukolsky equation can be used to find the eigenvectors in a more efficient way, as shown in [44].
Our numerical implementation uses the pseudo-spectral method, where all quantities are expanded in Chebyshev polynomials radially and spin-weighted spherical harmonics in the angular directions. Multiplications are performed using fast transforms from the spectral basis to a grid. All results in this paper have radial expansion truncated at degree $n_{\text{max}}=63$, the spherical harmonic expansion truncating at $\ell_{\text{max}}=31$ and with 256 bit floating point numbers from the MultiFloats library [45].
To obtain the source term $S_{4}^{(2)}$, we take a different approach from the analytic treatment used in Paper I—here, the source is computed numerically from the Einstein Tensor. To be specific, we first consider the second-order perturbation of the metric
$g_{ab}=g^{\circ}_{ab}+\epsilon h_{ab}^{(1)}+\epsilon^{2}h_{ab}^{(2)}+\mathcal{O}\left(\epsilon^{3}\right)\,,$ (3)
where $\epsilon$ is a book-keeping parameter that counts the orders. The Einstein Tensor, as a functional of the metric, can then be expanded as
$G_{ab}[g_{ab}]=$ $\epsilon G_{ab}^{(1)}[h_{ab}^{(1)}]+\epsilon^{2}G_{ab}^{(1)}[h_{ab}^{(2)}]+\epsilon^{2}G_{ab}^{(2)}[h_{ab}^{(1)},h_{ab}^{(1)}]$
$+\mathcal{O}\left(\epsilon^{3}\right)\,,$ (4)
where $G_{ab}^{(1)}$ and $2G_{ab}^{(2)}$ are the first and second functional derivatives of the Einstein Tensor with respect to the metric. To first order, we take $h_{ab}^{(1)}$ to be a sum of QNMs. We solve for $h_{ab}^{(1)}$ by numerically computing the Hertz potential for QNMs in the Outgoing Radiation Gauge (ORG), using the separation technique described in [44]. Note that for a QNM with frequency $\omega$, $h_{ab}^{(1)}$ will have both frequencies $\omega$ and $-\bar{\omega}$, which comes from the complex conjugations necessary to make $h_{ab}^{(1)}$ real.
Next, by imposing Einstein’s equations at second order, we get the equation
$G_{ab}^{(1)}[h_{ab}^{(2)}]=-G_{ab}^{(2)}[h_{ab}^{(1)},h_{ab}^{(1)}]\,,$ (5)
thus $h_{ab}^{(2)}$ satisfies the linearized Einstein equations with $-G_{ab}^{(2)}[h_{ab}^{(1)},h_{ab}^{(1)}]/8\pi$ acting as the effective stress-energy tensor at this order. After computing $G_{ab}^{(2)}[h_{ab}^{(1)},h_{ab}^{(1)}]$ from the linearized metric $h_{ab}^{(1)}$, we convert it to the source $S_{4}^{(2)}$ for $\Psi_{4}^{(2)}[h_{ab}^{(2)}]$ in Eq. (2). Note that while Eq. (5) only provides the source for the part of $\Psi_{4}$ that depends linearly on $h_{ab}^{(2)}$, in ORG quadratic contribution to $\Psi_{4}$ from $h_{ab}^{(1)}$ vanishes.
Finally, we numerically solve the Teukolsky equations with no-incoming boundary conditions to obtain $\Psi_{4}^{(2)}$ and $\Psi_{0}^{(2)}$ for the quadratic modes. More details can be found in the Supplementary Material.
III Classification of mode couplings
Quadratic coupling channels can be viewed in two ways. The first perspective traces the relationship from linear parent modes to quadratic daughter modes. Consider two linear QNMs $(\ell_{1},m_{1},n_{1},p_{1})$ and $(\ell_{2},m_{2},n_{2},p_{2})$, and amplitudes $A_{1}$ and $A_{2}$. As discussed above, the linearized metric $h_{ab}^{(1)}$ has four frequency components $\omega_{\ell_{1}m_{1}n_{1}}^{p_{1}}$, $-\bar{\omega}_{\ell_{1}m_{1}n_{1}}^{p_{1}}$, $\omega_{\ell_{2}m_{2}n_{2}}^{p_{2}}$, and $-\bar{\omega}_{\ell_{2}m_{2}n_{2}}^{p_{2}}$, with the amplitudes $A_{1}$, $\bar{A}_{1}$, $A_{2}$ and $\bar{A}_{2}$. Consequently, the source term in Eq. (5) yields four quadratic frequencies:
$\omega_{\ell_{1}m_{1}n_{1}}^{p_{1}}+\omega_{\ell_{2}m_{2}n_{2}}^{p_{2}},$ $\omega_{\ell_{1}m_{1}n_{1}}^{p_{1}}-\bar{\omega}_{\ell_{2}m_{2}n_{2}}^{p_{2}},$
$-\bar{\omega}_{\ell_{1}m_{1}n_{1}}^{p_{1}}+\omega_{\ell_{2}m_{2}n_{2}}^{p_{2}},$ $-\bar{\omega}_{\ell_{1}m_{1}n_{1}}^{p_{1}}-\bar{\omega}_{\ell_{2}m_{2}n_{2}}^{p_{2}},$
with source terms proportional to the products $A_{1}A_{2}$, $A_{1}\bar{A}_{2}$, $\bar{A}_{1}A_{2}$ and $\bar{A}_{1}\bar{A}_{2}$, respectively. This gives rise to four distinct channels through which a pair of linear QNMs can generate QQNMs at different frequencies.
The first perspective, which maps two parent modes to four daughter modes, is often conceptually inconvenient. Instead, it is more intuitive to adopt the opposite viewpoint: given a QQNM with a certain frequency, what are its parent modes? To formalize this, we will use the following notation to denote a QQNM:
$\begin{pmatrix}\ell_{1},&m_{1},&n_{1},&p_{1}\\ \ell_{2},&m_{2},&n_{2},&p_{2}\end{pmatrix}.$ (6)
Its time and azimuthal dependence is expressed as:
$\sim e^{-i(\omega_{\ell_{1},m_{1},n_{1}}^{p_{1}}+\omega_{\ell_{2},m_{2},n_{2}}^{p_{2}})t}e^{i(m_{1}+m_{2})\phi}.$ (7)
With this notation, several key observations can be made. First, the QQNM appears only in the $(m_{1}+m_{2})$ GW harmonics due to the azimuthal angular selection rule. Second, the notation exhausts all possible QQNMs. Third, different pairs of modes (up to exchange symmetry $1\leftrightarrow 2$) generally produce distinct frequencies. An exception occurs for Schwarzschild BHs, where spherical symmetry makes QNM frequencies independent of $m$. Nonetheless, different $(m_{1},m_{2})$ pairs indicate distinct coupling channels. The total QQNM amplitude in a GW harmonic is the sum of all possible pairs $(m_{1},m_{2})$ with $m_{1}+m_{2}$ fixed.
The QQNM in Eq. (6) is generated by four linear QNMs:
$(\ell_{1},m_{1},n_{1},p_{1}),$ $(\ell_{1},-m_{1},n_{1},-p_{1}),$
$(\ell_{2},m_{2},n_{2},p_{2}),$ $(\ell_{2},-m_{2},n_{2},-p_{2}),$
with respective amplitudes $A_{1}$, $A_{-1}$, $A_{2}$, and $A_{-2}$. It is worth noting that while the frequency values of $(\ell_{1},m_{1},n_{1},p_{1})$ and $(\ell_{1},-m_{1},n_{1},-p_{1})$ are related via Eq. (1), their amplitudes $A_{1}$ and $A_{-1}$ are independent, except in the case of nonprecessing systems, where because of the reflection symmetry we have
$A_{-1}=(-1)^{\ell_{1}}\bar{A}_{1}.$ (8)
The four linear QNMs are coupled through four channels
$(\ell_{1},m_{1},n_{1},p_{1})\otimes(\ell_{2},m_{2},n_{2},p_{2}),$ ++++,+ + , (9a)
$(\ell_{1},m_{1},n_{1},p_{1})\otimes(\ell_{2},-m_{2},n_{2},-p_{2}),$ ++-,+ - , (9b)
$(\ell_{1},-m_{1},n_{1},-p_{1})\otimes(\ell_{2},m_{2},n_{2},p_{2}),$ +-+,- + , (9c)
$(\ell_{1},-m_{1},n_{1},-p_{1})\otimes(\ell_{2},-m_{2},n_{2},-p_{2}),$--.- - . (9d)
Below we will refer to them as +++++ +, ++-+ -, +-+- +, and --- -, respectively. For example, the (+)(+-)( + - ) channel corresponds to the interaction where the linear mode $(\ell_{2},-m_{2},n_{2},-p_{2})$ generates a frequency component $-\bar{\omega}_{\ell_{2}-m_{2}n_{2}}^{-p_{2}}$ in the linearized metric, which is numerically equal to $\omega_{\ell_{2}m_{2}n_{2}}^{p_{2}}$ due to Eq. (1). This component then interacts with the frequency $\omega_{\ell_{1}m_{1}n_{1}}^{p_{1}}$ generated by the first linear mode $(\ell_{1},m_{1},n_{1},p_{1})$. Denoting the amplitude of the QQNM as $A_{Q}$, we have
$A_{Q}=$ $R_{++}A_{1}A_{2}+R_{+-}A_{1}\bar{A}_{-2}+R_{-+}\bar{A}_{-1}A_{2}$
$+R_{--}\bar{A}_{-1}\bar{A}_{-2}.$ (10)
where the coefficients $R$’s are fully determined by the dimensionless spin of a Kerr BH. Here we use spin-weighted spherical/spheroidal harmonics as the mode basis to facilitate direct comparison with previous results, such as those from numerical relativity and from Paper I. Alternatively, modes can also be classified by parity, see the Supplementary Material for more details.
For nonprecessing binaries, Eq. (10) reduces to
$R^{\rm total}\equiv\frac{A_{Q}}{A_{1}A_{2}}=$ $R_{++}+(-1)^{\ell_{2}}R_{+-}$
$+(-1)^{\ell_{1}}R_{-+}+(-1)^{\ell_{1}+\ell_{2}}R_{--},$ (11)
where we have used Eq. (8). Notice that the right-hand side (RHS) is an intrinsic property of a Kerr BH, independent of the linear mode amplitudes. We define the RHS as the total amplitude ratio $R^{\rm total}$, which characterizes the excitability of the QQNM.
The ()(--)( - - ) channel deserves more attention. We find that $R_{--}$ always vanishes identically, regardless of the types of parent modes and the spin of the Kerr BH. This result follows directly from the second-order Teukolsky equation in Eq. (2), where the source $S_{4}^{(2)}$ takes the following form in ORG [18]:
$S_{4}^{(2)}=(\ldots)^{(1)}\Psi_{4}^{(1)}+(\ldots)^{(1)}\Psi_{3}^{(1)}+(\ldots)^{(1)}\Psi_{2}^{(1)}.$ (12)
The structure of $S_{4}^{(2)}$ explicitly illustrates how two linear parent modes couple. The first parent mode contributes the three Weyl scalars $\Psi_{4}^{(1)}$, $\Psi_{3}^{(1)}$, and $\Psi_{2}^{(1)}$, which are then multiplied by the terms contributed by the second parent mode (indicated in parentheses). For a given linear QNM $(l,m,n,p)$, the expressions for $\Psi_{4}^{(1)}$, $\Psi_{3}^{(1)}$, and $\Psi_{2}^{(1)}$ were analytically derived in Paper I [18], see Eqs. (38), (42), and (46) therein. An intriguing feature emerges: all three Weyl scalars contain only a single frequency component, $\omega_{lmn}^{p}$, with no contribution from their mirrored counterpart, $-\bar{\omega}_{lmn}^{p}$. As a result, the source in Eq. (12) contains no terms resulting from the product of two mirrored frequency components.
IV Results and Comparisons
The QQNM $\begin{pmatrix}220+\\ 220+\end{pmatrix}$ in the $l=m=4$ harmonic has been extensively discussed in the literature [12, 10, 11, 13, 14, 19, 20, 21, 18]. In particular, in Paper I, we found that the computed amplitude ratio, which is essentially $R_{++}$ in Eq. (10), was smaller than those extracted from numerical relativity [13, 14]. We speculated that the difference was due to parity freedom in the linear parent modes (see Sec. VI. B in [18]). This was later confirmed in [19, 20, 21] for Schwarzschild BHs. In this work, after classifying all mode-coupling channels, we revisit the discrepancy for Kerr BHs.
Figure 1: A comparison between the complex contour method (cyan dots, from Paper I), time-domain numerical simulations (green, blue, and orange dots, from Zhu et al. [13]), and our current hyperboloidal slicing method (gray, black, and magenta curves). The gray curve represents the total amplitude ratio $R^{\rm total}$ for nonprecessing binaries [Eq. (11)]; the black curve corresponds to $R_{++}$ in Eq. (9a); and the magenta curve shows $R_{+-}$ in Eq. (9b). The mode amplitudes extracted by Zhu et al. [13] used spin-weighted spherical harmonics, while Paper I adopted spin-weighted spheroidal harmonics. Here we use spin-weighted spheroidal harmonics to make a fair comparison with Paper I. Results between these two angular bases differ by $\leq 1\%$ (see Supplementary Material).
As shown by the black solid lines in Fig. 1, we first compute the magnitude and phase of $R_{++}$ as functions of the dimensionless spin $\chi$. Good agreement is observed with the results obtained in Paper I, as represented by cyan dots. The magenta curves correspond to the results for $R_{+-}$. Unlike $R_{++}$, the magnitude of $R_{+-}$ exhibits a weaker and non-monotonic dependence on spin, and its value is smaller than that of $R_{++}$ by a factor of $\sim 7$.
For nonprecessing binaries, the total excitability of the QQNM is characterized by $R^{\rm total}$ [Eq. (11)]. The gray curves in Fig. 1 show its dependence on spin. For a Schwarzschild BH, $R^{\rm total}=0.1534e^{-0.06787i}$, in agreement with the values reported in [19, 20, 21]. For Kerr BHs, the results (for $R^{\rm total}$) extracted from numerical-relativity scattering experiments [13] are shown as green, blue, and orange dots in Fig. 1. Our findings are consistent with theirs.
We note that the values of $R$’s depend on the choice of angular basis — Paper I adopts spin-weighted spheroidal harmonics whereas numerical relativity simulations use spin-weighted spherical harmonics. In the Supplementary Material, we compare $R$ across these two bases, finding that they differ by only $\sim 1\%$ even when $\chi=0.95$. For a fair comparison with Paper I, we use spin-weighted spheroidal harmonics for the black and magenta curves in Fig. 1.
Besides $\begin{pmatrix}220+\\ 220+\end{pmatrix}$, we also study five more QQNMs, including $\begin{pmatrix}330+\\ 220+\end{pmatrix}$, $\begin{pmatrix}440+\\ 2\text{-}20-\end{pmatrix}$, $\begin{pmatrix}220+\\ 200+\end{pmatrix}$, $\begin{pmatrix}220+\\ 200-\end{pmatrix}$, and $\begin{pmatrix}330+\\ 2\text{-}20-\end{pmatrix}$. In particular, we use polynomials to fit their $R$ for different channels as functions of $\chi$:
$R=Ae^{i\phi},\quad\text{with}~{}A=\sum_{i=0}^{7}A_{i}\chi^{i},\phi=\sum_{i=0}^{7}\phi_{i}\chi^{i}.$ (13)
The values of the coefficients are provided in tables in the Supplementary Material.
V Forecast on detectability
Using the total amplitude ratio $R^{\rm total}$ from Eq. (11), we can estimate the SNR of various QQNMs emitted by nonprecessing binaries. To be specific, for a given binary system, we first adopt the recent surrogate model postmerger [8] to compute linear mode amplitudes, along with the mass and spin of the remnant BH. We then use our fitting formula for $R^{\rm total}$ in Eq. (13) to estimate the amplitude of the corresponding QQNMs. Finally, we follow the method outlined in [46, 47, 42] to compute their SNRs.
In the following, we focus on a 40 km CE [48] and LISA [49]. For simplicity, we average the SNRs over all angles, which yields the orthogonal condition between two antenna patterns $F_{+,\times}$ for CE:
$F_{+,\times}^{2}\right>=\frac{1}{5},\quad\left<F_{+}F_{\times}\right>=0.$
For LISA, there is an additional factor of $3/2$. As pointed out in Paper I, a QQNM is present in different $\ell$ harmonics. Its total SNR is given by given by
$\rho=\sqrt{\sum_{\ell}\rho_{\ell,m}^{2}}.$ (14)
In practice, we find that the first $\ell$ harmonic typically dominates the total SNR.
(a) $\begin{pmatrix}3,&3,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$
(b) $\begin{pmatrix}3,&3,&0,&+\\ 2,&-2,&0,&-\end{pmatrix}$
Figure 2: The SNR dependence of two QQNMs, $\begin{pmatrix}3,&3,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$, and $\begin{pmatrix}3,&3,&0,&+\\ 2,&-2,&0,&-\end{pmatrix}$, on progenitor binaries’ mass ratio $q$ and effective spin $\chi_{\rm eff}$. The systems have a total mass of $100M_{\odot}$ and a luminosity distance of 410 Mpc. CE sensitivity is used.
In this paper, we adopt a conservative estimate by choosing the starting time to be $10M_{\rm tot}$ after the peak, where $M_{\rm tot}$ is the total mass of the binary system. Giesler et al. [50] demonstrate stable QNM fits at earlier times. Selecting an earlier starting time can moderately increase the SNRs, since QQNMs decay faster than linear fundamental modes.
The SNR of $\begin{pmatrix}2,&2,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$ was investigated in [42]. In the Supplementary Material, we present our results. A major difference between us is in the SNR’s dependence on the mass ratio $q$. In their Fig. 6, the SNR peaks at $q\gtrsim 2$, while our more accurate calculation of the amplitude ratio shows a monotonic decrease with $q$. Therefore, the golden binary for detecting the QQNM is an equal-mass system with maximally positive spins.
The QQNM $\begin{pmatrix}3,&3,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$ was also identified in numerical relativity waveforms. In Fig. 2a, we present its SNR dependence on the mass ratio $q$ and the effective inspiral spin $\chi_{\rm eff}$. For simplicity, we assume both component spins $\chi_{1,2}$ are equal, namely $\chi_{1}=\chi_{2}=\chi_{\rm eff}$. The optimal binary configuration is $q\sim 2.75$ with maximally positive $\chi_{\rm eff}$. By setting the total binary mass $M_{\rm tot}$ to $100M_{\odot}$, and the luminosity distance $D_{L}$ to 410 Mpc, the SNR detected by CE can exceed 8 for some parameter space. With LISA, the SNR dependence on $q$ and $\chi_{\rm eff}$ remains the same. Assuming $M_{\rm tot}=8\times 10^{6}M_{\odot}$ and a redshift of 1, we find its SNR ranges from 16 to 653. The corresponding plot is available in the Supplementary Material.
The QQNM $\begin{pmatrix}3,&3,&0,&+\\ 2,&-2,&0,&-\end{pmatrix}$ is weaker. Comparing to $\begin{pmatrix}3,&3,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$, this QQNM has a smaller frequency and appears in the $(\ell\geq 2,m=1)$ harmonics. As shown in Fig. 2b, its SNR still peaks at $q\sim 2.75$ with maximally positive $\chi_{\rm eff}$. Using the same binaries, the SNR ranges from $0.1$ to $10.6$ for CE, and from $0.5$ to $59.8$ for LISA (see the LISA plot in the Supplementary Material).
Here we focus on two typical binary systems for CE and LISA by fixing their $M_{\rm tot}$ and $D_{L}$. The impact of $M_{\rm tot}$ on SNR is two-fold: (1) it scales the overall strain linearly; and (2) it sets the central frequency of a QQNM. In the Supplementary Material, we show how the SNR of $\begin{pmatrix}3,&3,&0,&+\\ 2,&-2,&0,&-\end{pmatrix}$ changes with $M_{\rm tot}$. For CE, SNR increases monotonically from 0.6 to 30 over $M_{\rm tot}\in[20,200]M_{\odot}$, while for LISA, it reaches a maximum for $M_{\rm tot}=5.3\times 10^{6}M_{\odot}$, with a value of 64.
VI Discussion
Together with Paper I, we have presented two independent frequency-domain methods for analyzing QQNMs — the complex contour technique and the hyperboloidal slicing approach. Both methods produce consistent results for $R_{++}$ of $\begin{pmatrix}220+\\ 220+\end{pmatrix}$. We have also classified the four possible coupling channels for a QQNM. Notably, the coupling between two mirrored modes (labeled by --- -) vanishes exactly. After accounting for all coupling channels in systems with reflection symmetry, our results for Kerr BHs match those from numerical simulations and time-domain fits.
We have performed a SNR survey on various QQNMs for nonprecessing binaries. Besides the known mode $\begin{pmatrix}220+\\ 220+\end{pmatrix}$, we considered two more: $\begin{pmatrix}3,&3,&0,&+\\ 2,&-2,&0,&-\end{pmatrix}$ and $\begin{pmatrix}3,&3,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$. For an event with $M_{\rm tot}=100M_{\odot}$ and $D_{L}=410$ Mpc, the SNR for CE may reach $\mathcal{O}(8)$ for binaries with $\chi_{\rm eff}>0.6$ and $q\in(2,5)$. For SgrA*-like binaries ($M_{\rm tot}=8\times 10^{6}M_{\odot}$) at a redshift of 1, LISA’s SNR may reach $\mathcal{O}(10^{2})$. We expect LISA to detect additional QQNMs. A more detailed survey is left for future work. Future studies could also explore precessing binaries, where QQNM excitation may encode richer information about linear parent modes and progenitors’ parity. It would be interesting to explore how to decipher the information from QQNM amplitudes.
Although this work, together with Paper I, has offered a comprehensive discussion on QQNMs, a complete account of all second-order effects has yet to be provided. Intriguing avenues for future study include the memory effect of a QNM, quadratic tails, and the possibility of resolving the nonlinear tails from numerical simulations, using techniques such as Cauchy-characteristic matching [51].
Acknowledgements.
VII Acknowledgments
We thank Hengrui Zhu for sharing the data in Fig. 1. We thank Andrew Spiers for pointing out the difference between the Campanelli–Lousto–Teukolsky equation and the reduced second-order Teukolsky equation. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. N.K. is supported by the Natural Science and Engineering Council of Canada.