### v3

# Quadratic Mode Couplings in Rotating Black Holes and Their Detectability

Abstract
Quadratic quasinormal modes encode fundamental properties of black hole spacetimes. They are also one of the key ingredients of nonlinearities of General Relativity in the ringdown stage of binary black hole coalescence. In this work, we classify all possible quadratic coupling channels of quasinormal modes for a generic Kerr black hole, and use a frequency-domain pseudospectral code with hyperboloidal slicing to calculate these couplings. After accounting for all the channels in systems with reflection symmetry, our results become consistent with those extracted from numerical simulations and time-domain fits. This agreement provides a compelling example demonstrating the success of black hole second-order perturbation theory. We also explore potential applications of our calculations in future ringdown data analysis by carrying out a detectability survey for various quadratic modes. We find that a few of them are observationally relevant for third-generation ground-based detectors like Cosmic Explorer, as well as the space-borne detector LISA (Laser Interferometer Space Antenna).
I Introduction
Black hole (BH) ringdown after binary BH coalescence encodes key information about the spacetime of the final BH [1, 2, 3], as well as the binary’s inspiral properties, such as the mass ratio, spins, etc [4, 5, 6, 7, 8, 9]. Theoretically, it is important to characterize various linear and non-linear signals in the ringdown stage and their connection to the inspiral parameters, as predicted by General Relativity (GR). In particular, understanding nonlinear signals not only enables direct tests of nonlinearities in GR, but also allows more theoretically meaningful ringdown fits at earlier times. This is particularly important for data analysis, as fitting the ringdown from earlier times generally leads to a higher signal-to-noise ratio (SNR).
Among various BH nonlinearities [10, 11, 12, 13, 14, 7, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], quadratic quasinormal modes (QQNMs) stand out due to their conceptual simplicity and their potential observability using current ringdown fitting techniques [27, 28, 29, 30, 31, 32, 33, 34]. QQNMs have been identified by fitting numerical time-domain simulation data at infinity [10, 11, 12, 13, 14, 7] and on the BH horizon [17]. However, a complete and accurate characterization of the modes relies on frequency domain methods, which is the focus of this work.
Along with an earlier work [18] (hereafter Paper I), we develop two independent frequency-domain methods to handle QQNMs for generic Kerr BHs. In Paper I, we analytically reconstructed the metric and solved the frequency-domain second-order Campanelli–Lousto-Teukolsky equation [35] along a complex contour. Alternatively, in this paper, we adopt the reduced second-order Teukolsky equation [36, 37] and use the pseudo-spectral method to numerically assemble its source term. Subsequently, we solve the equation spectrally in hyperboloidal slicing [38, 39, 40]. Both methods generate consistent results for the coupling coefficient of the prograde mode $\ell=m=2,n=0$ for BH spins $\chi\in[0,0.95]$. Furthermore, we extend our discussions in Paper I by providing a comprehensive classification of all possible mode-coupling channels that contribute to the same frequency of quadratic modes. This classification allows us to resolve the disagreement between our previous QQNM calculations in Paper I and numerical simulations.
Leveraging the recent ringdown surrogate model [^1] See also [9]. [8], we discuss potential applications of our calculations in future ringdown data analysis. As an example, we carry out a parameter survey to forecast the SNRs of several quadratic modes measured by Cosmic Explorer (CE) and Laser Interferometer Space Antenna (LISA) [41, 42], complementing the recent study by Yi et al. [43]. Our findings indicate that multiple QQNMs are relevant for future ringdown analysis.
Throughout this work we adopt the geometric unit by setting $G=c=1$. Complex conjugates are represented by overlines. We use $(l,m,n,p)$ to stand for a QNM, where $p={\rm sgn}({\rm Re}(\omega))$; and $\omega$ to denote frequency values. Due to the reflection symmetry of Kerr, QNMs come in pairs with frequency values related via
$\omega^{p}_{\ell mn}=-\bar{\omega}^{-p}_{\ell-mn},$ (1)
where modes with $(m,p)$ and $(-m,-p)$ (and the same $l,n$) are mirrors of each other.
II QQNMs with hyperboloidal slicing
QNMs are known to diverge at spatial infinity and at the bifurcate horizon. While this divergence poses no physical issue because they do not lie in the region where the QNMs are a good approximation, it can cause numerical challenges. In Paper I, we deal with the divergence by using the complex plane approach. Here, we adopt an alternative method—hyperboloidal time slicing—which avoids the problematic regions. This enables direct numerical treatments of the reduced second-order Teukolsky equation [36, 37]
$\mathcal{T}\left[\Psi_{4}^{(2)}\right]=S_{4}^{(2)}.$ (2)
Using the hyperboloidal slicing, it has been shown that QNMs are eigenvectors of the non-normal time evolution operator [44]. The separability of the Teukolsky equation can be used to find the eigenvectors in a more efficient way, as shown in [45].
To obtain the source term $S_{4}^{(2)}$, we take a different approach from the analytic treatment used in Paper I—here, the source is computed numerically from the Einstein Tensor. Specifically, we consider the second-order perturbation of the metric
$g_{ab}=g^{\circ}_{ab}+\epsilon h_{ab}^{(1)}+\epsilon^{2}h_{ab}^{(2)}+\mathcal{O}\left(\epsilon^{3}\right)\,,$ (3)
where $\epsilon$ is a book-keeping parameter that counts the orders. The Einstein Tensor, as a functional of the metric, can then be expanded as
$G_{ab}[g_{ab}]=$ $\epsilon G_{ab}^{(1)}[h_{ab}^{(1)}]+\epsilon^{2}G_{ab}^{(1)}[h_{ab}^{(2)}]+\epsilon^{2}G_{ab}^{(2)}[h_{ab}^{(1)},h_{ab}^{(1)}]$
$+\mathcal{O}\left(\epsilon^{3}\right)\,,$ (4)
where $G_{ab}^{(1)}$ and $2G_{ab}^{(2)}$ are the first and second functional derivatives of the Einstein Tensor with respect to the metric. To first order, we take the GW strain to be a sum of QNMs. We solve for $h_{ab}^{(1)}$ by numerically computing the Hertz potential for QNMs in the Outgoing Radiation Gauge (ORG), using the separation technique described in [45]. Note that for a QNM with frequency $\omega$, $h_{ab}^{(1)}$ will have both frequencies $\omega$ and $-\bar{\omega}$, which comes from the complex conjugations necessary to make $h_{ab}^{(1)}$ real.
Next, by imposing Einstein’s equations at second order, we get the equation
$G_{ab}^{(1)}[h_{ab}^{(2)}]=-G_{ab}^{(2)}[h_{ab}^{(1)},h_{ab}^{(1)}]\,,$ (5)
thus $h_{ab}^{(2)}$ satisfies the linearized Einstein equations with $-G_{ab}^{(2)}[h_{ab}^{(1)},h_{ab}^{(1)}]/8\pi$ acting as the effective stress-energy tensor at this order. After computing $G_{ab}^{(2)}[h_{ab}^{(1)},h_{ab}^{(1)}]$ from the linearized metric $h_{ab}^{(1)}$, we convert it to the source $S_{4}^{(2)}$ for $\Psi_{4}^{(2)}[h_{ab}^{(2)}]$ in Eq. (2). Note that while Eq. (5) only provides the source for the part of $\Psi_{4}$ that depends linearly on $h_{ab}^{(2)}$, in ORG quadratic contribution to $\Psi_{4}^{(2)}$ from $h_{ab}^{(1)}$ vanishes.
Finally, we numerically solve the Teukolsky equations with no-incoming boundary conditions to obtain $\Psi_{4}^{(2)}$ and $\Psi_{0}^{(2)}$ for the quadratic modes. More details can be found in the Supplementary Material.
III Classification of mode couplings
Quadratic coupling channels can be viewed in two perspectives. The first traces the relationship from linear parent modes to quadratic daughter modes. Consider two linear QNMs $(\ell_{1},m_{1},n_{1},p_{1})$ and $(\ell_{2},m_{2},n_{2},p_{2})$, with strain amplitudes $A_{1}$ and $A_{2}$. The corresponding linearized metric $h_{ab}^{(1)}$ has four frequency components $\omega_{\ell_{1}m_{1}n_{1}}^{p_{1}}$, $-\bar{\omega}_{\ell_{1}m_{1}n_{1}}^{p_{1}}$, $\omega_{\ell_{2}m_{2}n_{2}}^{p_{2}}$, and $-\bar{\omega}_{\ell_{2}m_{2}n_{2}}^{p_{2}}$, with the amplitudes $A_{1}$, $\bar{A}_{1}$, $A_{2}$ and $\bar{A}_{2}$, generated by the frequency pairs of the linear modes $(\ell_{1},m_{1},n_{1},p_{1})$ and $(\ell_{2},m_{2},n_{2},p_{2})$. Consequently, the source term in Eq. (5) yields four quadratic frequencies:
$\omega_{\ell_{1}m_{1}n_{1}}^{p_{1}}+\omega_{\ell_{2}m_{2}n_{2}}^{p_{2}},$ $\omega_{\ell_{1}m_{1}n_{1}}^{p_{1}}-\bar{\omega}_{\ell_{2}m_{2}n_{2}}^{p_{2}},$
$-\bar{\omega}_{\ell_{1}m_{1}n_{1}}^{p_{1}}+\omega_{\ell_{2}m_{2}n_{2}}^{p_{2}},$ $-\bar{\omega}_{\ell_{1}m_{1}n_{1}}^{p_{1}}-\bar{\omega}_{\ell_{2}m_{2}n_{2}}^{p_{2}},$
with source terms proportional to the products $A_{1}A_{2}$, $A_{1}\bar{A}_{2}$, $\bar{A}_{1}A_{2}$ and $\bar{A}_{1}\bar{A}_{2}$, respectively. This gives rise to four distinct channels through which two linear QNMs can generate QQNMs at different frequencies.
The first perspective, which maps two parent modes to four daughter modes, is often practically inconvenient. Instead, it is more intuitive to adopt the opposite viewpoint: given a certain QQNM frequency, what are its parent modes? To formalize this, we will use the following notation to denote a QQNM frequency component:
$\begin{pmatrix}\ell_{1},&m_{1},&n_{1},&p_{1}\\ \ell_{2},&m_{2},&n_{2},&p_{2}\end{pmatrix}.$ (6)
Its time and azimuthal dependence is expressed as:
$\sim e^{-i(\omega_{\ell_{1},m_{1},n_{1}}^{p_{1}}+\omega_{\ell_{2},m_{2},n_{2}}^{p_{2}})t}e^{i(m_{1}+m_{2})\phi}.$ (7)
This notation represents the set of all QQNMs with that frequency, including various channels, and is not restricted to fixed daughter spherical/spheroidal harmonic. With this notation, several key observations can be made. First, QQNMs with this frequency only appear in the $(m_{1}+m_{2})$ GW harmonics due to the azimuthal angular selection rule. Second, different pairs of modes (up to exchange symmetry $1\leftrightarrow 2$) generally produce distinct frequencies. An exception occurs for Schwarzschild BHs, where spherical symmetry makes QNM frequencies independent of $m$. Nonetheless, different $(m_{1},m_{2})$ pairs indicate distinct coupling channels. The total QQNM amplitude in a GW harmonic is the sum of all possible pairs $(m_{1},m_{2})$ with $m_{1}+m_{2}$ fixed.
The QQNM frequency in Eq. (6) is generated by four linear QNMs:
$(\ell_{1},m_{1},n_{1},p_{1}),$ $(\ell_{1},-m_{1},n_{1},-p_{1}),$
$(\ell_{2},m_{2},n_{2},p_{2}),$ $(\ell_{2},-m_{2},n_{2},-p_{2}),$
with respective strain amplitudes $A_{1}$, $A_{-1}$, $A_{2}$, and $A_{-2}$. It is worth noting that while the frequency values of $(\ell_{1},m_{1},n_{1},p_{1})$ and $(\ell_{1},-m_{1},n_{1},-p_{1})$ are related via Eq. (1), their amplitudes $A_{\pm 1}$ are independent, except in nonprecessing systems, where reflection symmetry imposes
$A_{-1}=(-1)^{\ell_{1}}\bar{A}_{1}.$ (8)
The four linear QNMs, coupled through the following four channels give rise to the same QQNM frequency:
$(\ell_{1},m_{1},n_{1},p_{1})\otimes(\ell_{2},m_{2},n_{2},p_{2}),$ ++++,+ + , (9a)
$(\ell_{1},m_{1},n_{1},p_{1})\otimes(\ell_{2},-m_{2},n_{2},-p_{2}),$ ++-,+ - , (9b)
$(\ell_{1},-m_{1},n_{1},-p_{1})\otimes(\ell_{2},m_{2},n_{2},p_{2}),$ +-+,- + , (9c)
$(\ell_{1},-m_{1},n_{1},-p_{1})\otimes(\ell_{2},-m_{2},n_{2},-p_{2}),$--.- - . (9d)
Below we will refer to them as +++++ +, ++-+ -, +-+- +, and --- -, respectively. For example, the (+)(+-)( + - ) channel corresponds to the interaction where the linear mode $(\ell_{2},-m_{2},n_{2},-p_{2})$ generates a frequency component $-\bar{\omega}_{\ell_{2}-m_{2}n_{2}}^{-p_{2}}$ in the linearized metric, which is numerically equal to $\omega_{\ell_{2}m_{2}n_{2}}^{p_{2}}$ due to Eq. (1). This component then interacts with the frequency $\omega_{\ell_{1}m_{1}n_{1}}^{p_{1}}$ generated by the first linear mode $(\ell_{1},m_{1},n_{1},p_{1})$. Denoting the strain amplitude of the QQNM as $A_{Q}$ (using either the spherical or spheroidal basis), we have
$A_{Q}=$ $R_{++}A_{1}A_{2}+R_{+-}A_{1}\bar{A}_{-2}+R_{-+}\bar{A}_{-1}A_{2}$
$+R_{--}\bar{A}_{-1}\bar{A}_{-2}.$ (10)
where the coefficients $R$’s are fully determined by the dimensionless spin of a Kerr BH, and the angular mode under consideration. Here we use spin-weighted spherical/spheroidal harmonics as the mode basis to facilitate direct comparison with previous results, such as those from numerical relativity and from Paper I. Alternatively, modes can also be classified by parity, see the Supplementary Material for more details.
For nonprecessing binaries, Eq. (10) reduces to
$R^{\rm total}\equiv\frac{A_{Q}}{A_{1}A_{2}}=$ $R_{++}+(-1)^{\ell_{2}}R_{+-}$
$+(-1)^{\ell_{1}}R_{-+}+(-1)^{\ell_{1}+\ell_{2}}R_{--},$ (11)
where we have used Eq. (8). Notice that the right-hand side (RHS) is an intrinsic property of a Kerr BH, independent of the linear mode amplitudes. We define the RHS as the total amplitude ratio $R^{\rm total}$, which characterizes the excitability of the QQNM.
The ()(--)( - - ) channel deserves more attention. We find that $R_{--}$ always vanishes identically, regardless of the types of parent modes and the spin of the Kerr BH. This result follows directly from the second-order Teukolsky equation in Eq. (2), where the source $S_{4}^{(2)}$ takes the following form in ORG [18]:
$S_{4}^{(2)}=(\ldots)^{(1)}\Psi_{4}^{(1)}+(\ldots)^{(1)}\Psi_{3}^{(1)}+(\ldots)^{(1)}\Psi_{2}^{(1)}.$ (12)
The structure of $S_{4}^{(2)}$ explicitly illustrates how two parent modes couple. The first parent mode contributes the three Weyl scalars $\Psi_{4}^{(1)}$, $\Psi_{3}^{(1)}$, and $\Psi_{2}^{(1)}$, which are then multiplied by the terms contributed by the second parent mode (indicated in parentheses). For a given linear QNM $(l,m,n,p)$, the expressions for $\Psi_{4}^{(1)}$, $\Psi_{3}^{(1)}$, and $\Psi_{2}^{(1)}$ were analytically derived in Paper I [18], see Eqs. (38), (42), and (46) therein. An intriguing feature emerges: all three Weyl scalars contain only a single frequency component, $\omega_{lmn}^{p}$, with no contribution from the $-\bar{\omega}_{lmn}^{p}$ frequency component in the linearized metric that is associated with the mode $(\ell,m,n,p)$. As a result, the source in Eq. (12) contains no terms resulting from the product of two conjugated frequency components.
IV Results and Comparisons
The QQNM $\begin{pmatrix}220+\\ 220+\end{pmatrix}$ in the $l=m=4$ harmonic has been extensively discussed in the literature [12, 10, 11, 13, 14, 19, 20, 21, 18]. In particular, in Paper I, we found that the computed amplitude ratio, $R_{++}$ for the $(4,4)$ mode in Eq. (10), was smaller than those extracted from numerical relativity [13, 14]. We speculated that the difference was due to parity freedom in the parent modes (see Sec. VI. B in [18]). This was later confirmed in [19, 20, 21] for Schwarzschild BHs. In this work, after classifying all mode-coupling channels, we revisit the discrepancy for Kerr BHs.
Figure 1: A comparison between the complex contour method (cyan dots, from Paper I), time-domain numerical simulations (green, blue, and orange dots, from Zhu et al. [13]), and our current hyperboloidal slicing method (gray, black, and magenta curves). The gray curve represents the total amplitude ratio $R^{\rm total}$ for nonprecessing binaries [Eq. (11)]; the black curve corresponds to $R_{++}$ in Eq. (9a); and the magenta curve shows $R_{+-}$ in Eq. (9b). The mode amplitudes extracted by Zhu et al.  [13] used spin-weighted spherical harmonics, while Paper I adopted spin-weighted spheroidal harmonics. Here we use spin-weighted spheroidal harmonics to make a fair comparison with Paper I. Results between these two angular bases differ by $\leq 1\%$ (see Supplementary Material).
As shown by the black solid lines in Fig. 1, we first compute the magnitude and phase of $R_{++}$ as functions of the dimensionless spin $\chi$. Good agreement is observed with the results obtained in Paper I, as represented by cyan dots. The magenta curves correspond to the results for $R_{+-}$. Unlike $R_{++}$, the magnitude of $R_{+-}$ exhibits a weaker and non-monotonic dependence on spin, and its value is smaller than that of $R_{++}$ by a factor of $\sim 7$.
For nonprecessing binaries, the total excitability of the QQNM is characterized by $R^{\rm total}$ [Eq. (11)]. The gray curves in Fig. 1 show its dependence on spin. For a Schwarzschild BH, $R^{\rm total}=0.1534e^{-0.06787i}$, in agreement with the values reported in [19, 20, 21]. For Kerr BHs, the results (for $R^{\rm total}$) extracted from numerical-relativity scattering experiments [13] are shown as green, blue, and orange dots in Fig. 1. Our findings are consistent with theirs.
We note that the values of $R$’s depend on the choice of angular basis — Paper I adopts spin-weighted spheroidal harmonics whereas numerical relativity simulations use spin-weighted spherical harmonics. In the Supplementary Material, we compare $R$ across these two bases, finding that they differ by only $\sim 1\%$ even when $\chi=0.95$. For a fair comparison with Paper I, we use spin-weighted spheroidal harmonics for the black and magenta curves in Fig. 1.
Besides $\begin{pmatrix}220+\\ 220+\end{pmatrix}$, we also study six more QQNMs, including $\begin{pmatrix}330+\\ 220+\end{pmatrix}$, $\begin{pmatrix}440+\\ 2\text{-}20-\end{pmatrix}$, $\begin{pmatrix}440+\\ 3\text{-}20-\end{pmatrix}$, $\begin{pmatrix}220+\\ 200+\end{pmatrix}$, $\begin{pmatrix}220+\\ 200-\end{pmatrix}$, and $\begin{pmatrix}330+\\ 2\text{-}20-\end{pmatrix}$. These modes are heuristically picked, motivated by the observational relevance of the $(3,3,0,+)$ mode [^2] Other linear modes, such as $(210+)$ [29, 46], are also relevant. and the the recent QQNMs discovered in numerical waveforms [47]. We use polynomials to fit their $R$ for different channels as functions of $\chi$:
$R=Ae^{i\phi},\quad\text{with}~{}A=\sum_{i=0}^{7}A_{i}\chi^{i},\phi=\sum_{i=0}^{7}\phi_{i}\chi^{i}.$ (13)
The values of the coefficients are provided in tables in the Supplementary Material.
V Implications for future ringdown data analysis
We now analyze the significance of the QQNMs on future ringdown data analysis by including all channels. For a binary system, the remnant surrogate model from [48] predicts the remnant’s spin, which can be used in Eq. (13) to compute quadratic amplitude ratios. When combined with the postmerger ringdown model [8], these ratios enable precise quadratic amplitude predictions. This information has the potential to improve future Bayesian analyses by narrowing prior ranges and reducing the parameter space. Note that our calculations include precessing systems.
Designing a systematic Bayesian framework that incorporates our results is beyond the scope of this paper. Instead, as a figure of merit, we illustrate a practical application of our calculations by estimating the SNR of various QQNMs that have not been previously discussed in the literature. Below, we follow the method outlined in [49, 50, 43] to compute these SNRs. For simplicity, we average the SNRs over all angles, which yields the orthogonal condition between two antenna patterns $F_{+,\times}$ for CE:
$F_{+,\times}^{2}\right>=\frac{1}{5},\quad\left<F_{+}F_{\times}\right>=0.$
For LISA, there is an additional factor of $3/2$. As pointed out in Paper I, a QQNM is present in different $\ell$ harmonics. Its total SNR is given by
$\rho=\sqrt{\sum_{\ell}\rho_{\ell,m}^{2}}.$ (14)
In practice, we find that the first $\ell$ harmonic typically dominates the total SNR.
Figure 2: The SNR dependence of the QQNM, $\begin{pmatrix}3,&3,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$, on progenitor binaries’ mass ratio $q$ and effective spin $\chi_{\rm eff}$, assuming CE. The binary systems have a total mass of ${60}M_{\odot}$ and a luminosity distance of 410 Mpc.
In this paper, we adopt a conservative estimate by choosing the starting time to be $10M_{\rm tot}$ after the peak, where $M_{\rm tot}$ is the total mass of the binary system. Selecting an earlier starting time can moderately increase the SNRs, since QQNMs decay faster than linear fundamental modes.
Figure 2 shows the SNR dependence of the QQNM $\begin{pmatrix}330+\\ 220+\end{pmatrix}$ on the mass ratio $q$ and the effective inspiral spin $\chi_{\rm eff}$. This analysis is performed for CE using $M_{\rm tot}=60M_{\odot}$ and a luminosity distance of 410 Mpc. For simplicity, we set both component spins equal, $\chi_{1}=\chi_{2}=\chi_{\rm eff}$. We can see that the optimal binary configuration occurs at $q\sim 2.75$ and maximally positive $\chi_{\rm eff}$, yielding SNR $\sim 4.7$.
Here, we focus on a representative binary by fixing $M_{\rm tot}$ and $D_{L}$. The total mass affects the SNR in two ways: (1) it scales the overall strain linearly; and (2) it sets the central frequency of a QQNM. Varying $M_{\rm tot}\in[20,200]M_{\odot}$, we find that SNR increases monotonically from 0.2 to 59. The corresponding plot is included in the Supplementary Material.
Besides $\begin{pmatrix}330+\\ 220+\end{pmatrix}$, Table 1 summarizes the SNRs of various QQNMs. Most are significant enough for LISA detection, while for CE, the SNRs fall just below the typical threshold ($\sim 8$) at the given $M_{\rm tot}$ and $D_{L}$, though they could be higher for more massive or closer binaries. Nevertheless, our results suggest the importance of subdominant QQNMs and the difficulty of future ringdown data analysis: While the QQNMs may lack sufficient SNR for precise parameter estimation, omitting them — as noted in [47] — can introduce systematic errors in linear mode fitting, destabilizing extracted amplitudes. Thus, a robust ringdown analysis pipeline will be essential for future detectors.
Table 1: Optimal SNRs of various QQNMs, assuming CE with $M_{\rm tot}=60M_{\odot}$ and $D_{L}=410$ Mpc, and LISA with $M_{\rm tot}=8\times 10^{6}M_{\odot}$ and a redshift of 1. The SNRs are obtained from optimal binary configurations that maximize SNR for the specified $M_{\rm tot}$ and $D_{L}$.
QQNM Optimal SNR
configuration CE LISA
$\begin{pmatrix}2,&2,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$ $q\sim 1$ large $\chi_{\rm eff}$ 9 970
$\begin{pmatrix}3,&3,&0,&+\\ 2,&2,&0,&+\end{pmatrix}$ $q\sim 2.75$ large $\chi_{\rm eff}$ 4.7 653
$\begin{pmatrix}4,&4,&0,&+\\ 2,&-2,&0,&-\end{pmatrix}$ $q\sim 4.25$ large $\chi_{\rm eff}$ 3 140
$\begin{pmatrix}3,&3,&0,&+\\ 2,&-2,&0,&-\end{pmatrix}$ $q\sim 2.75$ large $\chi_{\rm eff}$ 5 60
$\begin{pmatrix}4,&4,&0,&+\\ 3,&-2,&0,&-\end{pmatrix}$ $q\sim 8$ large $\chi_{\rm eff}$ $0.084$ 2.1
VI Discussion
Together with Paper I, we have presented two independent frequency-domain methods for analyzing QQNMs — the complex contour technique and the hyperboloidal slicing approach. Both methods produce consistent results for $R_{++}$ of $\begin{pmatrix}220+\\ 220+\end{pmatrix}$. We have also classified the four possible coupling channels for a QQNM. Notably, the coupling between two mirrored modes (labeled by --- -) vanishes exactly. After accounting for all coupling channels in systems with reflection symmetry, our results for Kerr BHs match those from numerical simulations and time-domain fits, which therefore resolves the tension between Paper I and Zhu et al. [13].
The agreement between our theoretical predictions and numerical relativity results provides a compelling quantitative demonstration—one of only a few such examples (e.g. [51, 52])—of the success of BH second-order perturbation theory. Our theoretical analysis reveals that the quadratic ratios reported in [13, 14] are limited to nonprecessing binaries. By contrast, the excitation of QQNMs in precessing binaries is expected to encode richer information about parent modes and progenitors’ parity. This opens an avenue for future studies to investigate how QQNMs amplitudes can infer precession parameters.
We have also illustrated potential applications of our calculations in future ringdown data analysis. As a figure of merit, we performed an SNR survey. Our results indicate that multiple QQNMs are likely to play a nonnegligible role in BH spectroscopy with future detectors. Therefore, a robust ringdown data analysis framework will be required to account for a substantial number of linear modes and QQNMs, in order to avoid systematic biases and prevent false alarms of deviations from GR.
Although this work, together with Paper I, has offered a comprehensive discussion on QQNMs, a complete account of all second-order effects has yet to be provided. Intriguing avenues for future study include the memory effect of a QNM, quadratic tails, and the possibility of resolving the nonlinear tails from numerical simulations, using techniques such as Cauchy-characteristic matching [53].
Acknowledgements.
VII Acknowledgments
We thank Hengrui Zhu for sharing the data in Fig. 1. We thank Andrew Spiers for pointing out the difference between the Campanelli–Lousto–Teukolsky equation and the reduced second-order Teukolsky equation. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. N.K. is supported by the Natural Science and Engineering Council of Canada.